OpenStudy (rebeccapink6398):

What is the area of this circle? (Express your answer using π.) A. A = 25π ft2 B. A = 5π ft2 C. A = 100π ft2 D. A = 10π ft2

OpenStudy (rebeccapink6398):



OpenStudy (rebeccapink6398):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

10 is the diameter right?

jimthompson5910 (jim_thompson5910):

it looks it, if so, then the radius is 10/2 = 5 so r = 5

jimthompson5910 (jim_thompson5910):

Plug r = 5 into A = pi*r^2 and tell me what you get

OpenStudy (rebeccapink6398):

78.5??

OpenStudy (rebeccapink6398):

@jim_thompson5910 ???

OpenStudy (rebeccapink6398):

hello???? are you there @jim_thompson5910 ??

jimthompson5910 (jim_thompson5910):

leave pi as it is

jimthompson5910 (jim_thompson5910):

so 5^2 = 25 making the answer to be 25pi